# Does GPT Really Get It? A Hierarchical Scale to Quantify Human and AI’s Understanding of Algorithms

### v1

@@ 3 Methods
We conducted a survey on students of algorithms courses at a premier CS-teaching university. Each student was assigned either the Euclidean or Ford-Fulkerson Algorithm at random, and was asked to rate their own understanding of the algorithm on a six-point scale. Each survey consisted of five test questions to test their understanding. There were three versions of the survey for each algorithm, assigned at random. The questions are available in Appendix A.
@@
The questions are designed to map to increasing levels of understanding.
  1. 1.
    A trivial instance of the problem. If the student understands the input and output space, the problem can be answered without calculation.
  2. 2.
    An intermediate instance of the problem. This is answerable without a calculator for most undergraduates, but requires the student to run some internal algorithm to compute.
  3. 3.
    A coding problem. This problem requires the student to translate a part of the algorithm to code.
  4. 4.
    Either an example problem or an explanation problem. This question asks the student to provide an example to illustrate a concept or explain the algorithm to a specified audience.
  5. 5.
    A counter-example or extension problem. This problem asks the student to reason about modifications to the algorithm.
@@
Several LLM models were presented with the same surveys given to the human participants; each survey was started in a fresh chat session, and the five questions were presented in the same order (within the survey, previous questions and responses were included in the chat history). We also included a system prompt to prime the LLM with the algorithm in question and encourage conciseness in the responses.
To ensure the statistical stability of the results, the LLM was also queried using randomized versions of the survey. For evaluation questions, the values in the questions were assigned uniformly at random within a given range. For the flow questions, the graph structure was also varied slightly. For the code questions, we took an example code implementation of the algorithm, randomly masked a line or group of lines, and asked the LLM to fill in the missing part. Lastly, for the example, explanation, and extension questions, we automatically generated variants of the question using GPT-4o (in a separate chat session). These questions were manually inspected to ensure they matched our criteria, and questions which were not useful were discarded.
Three versions of GPT were tested: GPT 3.5 Turbo, GPT 4, and GPT-4o.
@@
The quality of explanations and summaries can be subjective; however, they offer a deep insight into the subject’s understanding of the material. We evaluate the correctness of explanations on three axes.
@@
    First, correctness; the explanation is accurate and includes the key ideas of the algorithm.
@@
    Second, audience adaption; the explanation is tuned to the knowledge and needs of the audience.
@@
    Third, intuitiveness; the explanation conveys intuition - for example via contrast, example, or analogy - and uses clear language.
Note that summaries and explanations are by definition selective, and are not necessarily complete (Mittelstadt, Russell, and Wachter 2019). In fact, the ability to identify the key ideas are part of what differentiates explanation (Level 4) from the production of instructions (Level 3). The explanation is awarded 2/3 of a point for each bullet, for a maximum score of 2 per question.
Hypotheses
Therefore, any student who answers a question at a later level is likely to be able to answer a question from an earlier level. As a consequence, the fraction of correct responses should be monotonically non-increasing with the level, and the conditional probability of a correct response for a question should be higher if the previous question was answered correctly.
2. The landscape of understanding for GPT will be different from that of humans.
Since explanations of the algorithms and their properties appear frequently in the training data, we expect GPT-4 to be successful at questions related to explanation and summary. For the same reason (exposure to large corpora of code in training data), we expect it to do well at coding questions. However, we expect it to struggle with execution, including Questions 1 and 2, which require calculation. We expect this to be true for to a greater extent for GPT-3.5, which appears to have weaker arithmetic abilities.
3. The performance of GPT-4 can be improved via techniques such as chain-of-thought prompting and the inclusion of example solutions. These are recommended practices (OpenAI 2024).
4 Results
Out of all the responses collected, 85% of students indicated that they understood the algorithm, with most students reporting that they “know the algorithm and have a fair understanding of it”. For further analysis, we only consider responses from students who stated that they understood the algorithm. Across all students, the accuracy on the questions was highest for Question 1, and decreased uniformly through Question 5 (Figure 2).
Among versions of GPT, the best performing version was GPT-4 (Figure 4). The response score of GPT-4 was close to that of graduate students, as shown in Figure 2 (although this statement has some caveats, discussed further under Qualitative Comparison). LLMs on average out-performed the undergraduate students, except on questions 1 and 2 (evaluation). We take a closer look at the score distribution in Figure 5. The LLM performance exhibited a similar accuracy curve to students, with the exception of question 3 - all three versions of GPT tested were able to reliably answer the coding question regardless of whether the previous problem was evaluated correctly. This trend is exhibited in Figure 6.
Figure 2: The average score per question across undergraduate and doctoral students who self-reported that they understood the algorithm. Number of records is $n=13$ and $n=10$ respectively. Also shown is the average score by GPT-4 across 57 randomized versions of the surveys.
Figure 3: The average score per question and algorithm across undergraduate and graduate students who self-reported that they understood the algorithm.
Figure 4: The average score between three versions of GPT, across 27 randomized surveys on Max Flow and 30 on GCD.
@@ 4 Results
Figure 6: Difference in score between questions at consecutive levels, for both students (orange) and versions of GPT (blue). The circles are the means and the bars are the standard deviations.
Alignment with Levels.
In Figure 7, we map the results back to the levels shown in Figure 1. We computed the average score across questions in each category for each algorithm. Both GPT 4 and GPT 4o demonstrated an increase in quality in execution (Levels 1 and 2), explanation (Level 4b), and extension (Level 5b). On the Ford-Fulkerson algorithm, it’s ability to produce examples improved only marginally, and no version of GPT was able to successfully answer the difficult counterexample questions.
(a) GPT 3.5 Turbo
(b) GPT 4
(c) GPT 4o
(d) Undergraduate ($n=5$)
(e) Doctoral ($n=5$)
(f) GPT 3.5 Turbo
(g) GPT 4
(h) GPT 4o
(i) Undergraduate ($n=8$)
(j) Doctoral ($n=5$)
Figure 7: Average score at each level on the Ford Fulkerson Algorithm (Top, a-e) and Euclidean Algorithm (Bottom, f-j) for each GPT version. Darker colors indicate higher average scores for that question category.
@@
Figure 8: Percent accuracy on 100 randomized max flow problems per GPT version. Shows both base accuracy (blue) and accuracy when a correct example was given in the chat history (orange).
We find that this strategy generally causes the response to mimic the structure of the example response. As shown in Figure 8, including an example response improves accuracy for GPT 3.5, but has little-to-no effect on GPT 4o, and marginally decreases the accuracy for GPT 4. One possible explanation for this phenomenon is that GPT 4 naturally responds to the prompt with effective chain of thought reasoning; therefore, instructing it to reason in a specific format does not improve its reasoning abilities, and may in fact interfere with them.
@@
This section will highlight patterns in responses that cannot be seen in the aggregate results.
Generating Examples. As shown in Figure 7, all versions of GPT struggled to produce an example of a graph satisfying a prescribed property. This could be attributed to GPT’s known difficulties with mathematical calculation - after all, in order to know if the execution of the algorithm satisfies a property, it must first execute the algorithm, which GPT generally cannot do without chain-of-thought prompting(Wei et al. 2022). However, our evidence suggests a deeper issue with categorizing inputs.
We asked the following question: “Give an example of a graph where the Ford-Fulkerson algorithm computes exactly six augmenting flows before terminating. Write the example as a list of edges and capacities.” This question is deliberately ambiguous, as the number of augmenting flows depends on the method used to find paths. This ambiguity, along with the relatively large number of paths, should steer the respondent away from a ‘trial and error’ method. The intended answer is a graph with a source s, a sink t, and six intermediate vertices connected to both, leading to six parallel paths from s to t.
In the human survey, out of nine students, six gave this exact answer (the other three did not attempt the question). It is reasonable to expect that these students did not mentally execute the Ford-Fulkerson algorithm in order to verify their answer. Instead, they likely had some knowledge of the concept of parallel paths, and were able to leverage this knowledge to retrieve an example.
@@
The Fibonacci Sequence. We asked two variant questions exploring the relationship between the Fibonacci sequence and the Euclidean algorithm. One question asked the respondent to explain why consecutive Fibonacci numbers are considered worst-case inputs for the Euclidean algorithm. A correct answer would point out that since $F(n+1)\text{mod}F(n)=F(n-1)$, the input size reduces by approximately $\phi=1.618\dots$ at each step, leading to the worst-case run time of $O(\log(a))$. The second question asked the respondent to find an integer $x$ such that computing $\text{gcd}(55,x)$ takes the maximum number of steps, and to explain their choice. This question did not include the word Fibonacci, and the two variants were never asked to the same student so as not to bias the result.
Out of the ten students asked the second question, four provided a correct response. Although the word Fibonacci was never stated, a few students seemed to intuit the motivation behind the question. One doctoral student responded:
It should be approximately 1/phi (golden ratio) times 55 in general I think. Since 55 is a Fibonacci number it is easy to guess that the longest sequence should be obtained by 34, the preceding Fibonacci number, since this results in the Euclidean algorithm going through the Fibonacci sequence down to 1.
Another stated,
55 is a fibonacci number so I believe the answer is x=34 although I am not sure. That would require you to take one step in each direction each time.
This question also highlighted a jump in performance from GPT 3.5 to GPT 4. GPT 3.5 got an average score of 0.5/2 on the question, while GPT 4 and 4o averaged 1.9 and 2.0 respectively. However, GPT 3.5 also seemed to ‘intuit’ that the question involved the Fibonacci sequence. In five out of twelve responses, GPT 3.5 responded with the correct answer $x=34$, but with an incomplete or incorrect explanation (correct answers with incomplete explanations were scored as 1.33/2, while correct answers with nonsensical explanations were scored as 0.67/2). One response read:
The number 34 in the interval 0 $<$ x $<$ 55 leads to the maximal number of recursive steps. This is because 55 and 34 are consecutive Fibonacci numbers, which tend to lead to the largest number of recursive steps in the Euclidean algorithm due to their relatively prime nature.
While it is true that consecutive Fibonacci numbers are relatively prime, this does not explain why it requires more recursive steps than other pairs of numbers (e.g., while 54 and 55 are relatively prime, computing GCD(54, 55) only takes two recursive steps). Other explanations given by GPT were completely incorrect:
The integer $x$ that results in the longest sequence of recursive calls when calculating GCD($55,x$) using the Euclidean algorithm is $x=34$.
Explanation: […]. In this case, since $34$ is the largest integer less than $55$ that has a common factor with $55$ (GCD(55, 34) = 1), it results in the longest sequence of recursive calls compared to other integers. This is because the smaller the common factor, the more iterations are needed to reach a remainder of 00.
Responses such as this support the theory that the production of a correct answer is not direct evidence that LLMs have understanding.
Trouble with Ropes. For both the algorithms, one question asked the respondent to explain the algorithm to a student struggling with certain key concepts. One tactic which emerged was to explain the algorithm using a physical analogy. Thirteen out of the twenty-eight LLM responses asked the student to imagine the Ford-Fulkerson algorithm using water flowing through a series of pipes, and ten out of twenty-five LLM responses described the Euclidean algorithm as a way to find the smallest piece of rope that equally divides two ropes or sticks of different lengths. In contrast, one out of five student responses described FF using a series of pipes, and none of the six student responses used a physical analogy to explain the Euclidean algorithm.
We found three out of ten responses from LLMs using the rope/stick analogy for the Euclidean algorithm included explanations which did not make physical sense. For example, one response from GPT 3.5 described the algorithm as follows:
Take the two rope lengths (let’s call them Rope A and Rope B), and see how many times the shorter rope (Ride (sic) A) can wrap around the length of the longer rope (Rope B) without hanging over.
It is obvious to a human reader that one cannot wrap the shorter rope around the longer. This matches findings that LLMs lack functional understanding of properties of real objects (Mahowald et al. 2024).
Hallucinations. Within correct responses, we found that large language models included nonsequitors and inaccurate statements. As mentioned above, one extension question asked why consecutive Fibonacci numbers are considered ‘worst case’ inputs for the Euclidean algorithm. GPT-4 got a score of 2/2 on this question 92% of the time. However, within a correct answer, it occasionally included incorrect statements. For example, in one response it included the following line (emphasis ours):
@@
1. **Structure of Fibonacci numbers**: Each Fibonacci number is the sum of the two preceding ones, with the sequence beginning as $F(0)=0$, $F(1)=1$, $F(2)=1$, $F(3)=2$, and so forth. This means each number in the sequence is relatively close to the sum of the two preceding numbers.
In another response, it stated that $\phi$ was “one of the most irrational numbers”. This highlights some need for caution when using GPT-4 as a teaching implement.
5 Discussion
We have presented a hierarchical scale for measuring understanding of algorithms in both humans and AI. In this section, we discuss implications, interpretations amd limitations of our approach and point to future directions.
Our results show a significant improvement between the GPT 3.5 and GPT 4 generations in the skills of evaluation, explanation, and counterfactual reasoning for the two tested algorithms.
All versions of GPT performed near-perfectly on the skill of representation. This is in line with other findings showing the competence of GPT in code generation (Vaithilingam, Zhang, and Glassman 2022; Savelka et al. 2023). On the flip side, this suggests that if an LLM is able to generate code, it may not necessarily be able to perform other reasoning tasks such as evaluation and explanation.
@@ 5 Discussion
Second, we note that in order to determine whether GPT is reasoning, the concept of reasoning itself needs to be precisely defined. What our study indicates is that GPT-4 has a sophisticated internal representation, e.g., its representation of the Euclidean algorithm includes its relationship to Linear Diophantine equations and mathematical properties of GCD. It is able to retrieve these properties under a variety of contexts and framings. In a similar vein, it is likely that a student answering these questions has also been exposed to these properties in their studies. When evaluating the usefulness of a representation, we do not necessarily need to account for how the representation was created.
@@
Our results do not imply that GPT-4 has, on average, a ‘better’ understanding of the tested algorithms than undergraduates. The student study was included as a basis of comparison, and to demonstrate that the hierarchy we present can also apply to humans. However, we work with a limited population size, and the difference is confounded by other factors such as subject fatigue. Further research is needed to compare the quality of GPT and human responses to questions pertaining to algorithms. We hope that our scale can be applied to other algorithmic and similarly precise realms of understanding.
@@ Appendix A Survey Questions
Figure 9: The six point-scale shown to students to self-report their understanding.
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm.
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm.
@@
    You are speaking with a mathematics student who understands modular arithmetic. Explain the proof that the Euclidean algorithm finds the greatest common denominator of two numbers.

### v3

@@ Methods
We conducted a survey on students of algorithms courses at a premier CS-teaching university. Each student was assigned either the Euclidean or Ford-Fulkerson Algorithm at random, and was asked to rate their own understanding of the algorithm on a six-point scale. Each survey consisted of five test questions to test their understanding. There were three versions of the survey for each algorithm, assigned at random. The questions are available in the Appendix.
@@
Several versions of ChatGPT were presented with the same surveys given to the human participants; each survey was started in a fresh chat session, and within the survey, previous questions and responses were included in the chat history. We also included a system prompt to prime GPT and encourage conciseness in the responses.
GPT was also queried using randomized versions of the survey. For evaluation questions, the input values were assigned uniformly at random within a given range. For the flow questions, the graph structure was also varied slightly. For the code questions, we took an example code implementation of the algorithm, masked a key section, and asked GPT to fill in the missing part. We also included several versions of the example, explanation, and extension questions.
@@
The quality of explanations and summaries can be subjective; however, they offer a deep insight into the subject’s understanding of the material. We evaluate the explanations on three axes.
@@
    Correctness; the explanation is accurate and includes the key ideas of the algorithm.
@@
    Audience adaptation; the explanation is tuned to the audience, and the level of detail matches their prior knowledge.
@@
    Intuitiveness; the explanation conveys intuition; via contrast, example, analogy etc. and uses clear language.
Summaries and explanations are by definition selective and not necessarily complete (Mittelstadt, Russell, and Wachter 2019). The ability to identify key ideas is part of what differentiates explanation (Level 4) from the production of instructions (Level 3). An explanation is awarded 2/3 of a point for each bullet, for a maximum score of 2 per question.
Results
Figure 2: The average scores across students who self-reported that they understood the algorithm. Number of records is $n=13$ (undergraduate) and $n=10$ (graduate) respectively. The average scores for GPT-4 are across 60 randomized versions of the surveys. Error bars are 95% confidence intervals.
Figure 3: The average score between three versions of GPT, across 30 random surveys for each of GCD and Max Flow. Error bars show the 95% confidence interval.
@@ Results
Hypothesis 1
The understanding hierarchy (Figure 1) captures depth of understanding.
Overall, 85% of students indicated they understood the algorithm, with most students reporting that they “know the algorithm and have a fair understanding of it”. For further analysis, we only consider students who stated that they understood the algorithm. We compare undergraduate and doctoral levels in Fig. 2. Across all students, the accuracy on the questions was highest for Question 1, and decreased uniformly through Question 5. Doctoral students performed better on average than undergraduates ($p<0.05$). They received higher scores on Q4 and Q5 ($p<0.05$), while the differences on Q1, Q2, and Q3 were not statistically significant.
Figure 5: The difference in mean performance between mathematical and natural language reasoning tasks on Ford-Fulkerson (Left) and the Euclidean algorithm (right). The top graphs show tasks at Level 4, while the bottom graphs show tasks at Level 5
Hypothesis 2
Newer generations of GPT understand algorithms at a higher level than older generations.
Among versions of GPT, GPT-4 and GPT-4o performed about the same, and the differences in their overall scores were not significant (Figure 3). Both GPT-4 and GPT-4o demonstrated an increase in score on every question compared to GPT-3.5 ($p<0.05$).
The response score of GPT-4 was close to that of graduate students, as shown in Figure 2. Doctoral students scored better than GPT-4 on the extension questions (Q5) to a statistically significant degree. LLMs on average out-performed the undergraduate students on questions 3, 4, and 5 ($p<0.05$), while the differences on Q1 and Q2 were not statistically significant.
Hypothesis 3
LLMs will exhibit a performance gap between natural language reasoning and mathematical reasoning tasks.
As shown in Figure 5, all three versions tested performed better on language tasks than on mathematical reasoning tasks for Ford-Fulkerson (significant with $p<0.05$) despite student performance being the same or slightly worse. For GCD, the versions performed better on language tasks than on mathematical reasoning tasks on Level 4, but the same or slightly worse on Level 5.
We also hypothesized that the performance on code tasks would be higher compared to the performance on evaluation and reasoning tasks. We find that this does hold. As shown in Figure 3, LLM performed better on the coding tasks (Q3) than on the evaluation tasks (Q2), while the students exhibited the opposite trend (Figure 2).
@@
Figure 6: Accuracy of GPT versions on 100 (randomized) trivial/intermediate max flow problems. Base accuracy is in blue and accuracy when prompted with a correct example is in orange. Error bars are 95% confidence intervals.
We find that this strategy generally causes the response to mimic the structure of the example response. As shown in Figure 6, including an example response improves accuracy for GPT-3.5, but has little to no effect on GPT-4o, and marginally decreases the accuracy for GPT-4. One possible explanation for this phenomenon is that GPT-4 naturally responds to the prompt with effective chain of thought reasoning; therefore, instructing it to reason in a specific format does not improve its reasoning abilities, and may in fact interfere with them.
@@
Generating Examples. As shown in Figure 5 (left), all versions of GPT struggled to produce an example of a graph satisfying a prescribed property. This could be attributed to GPT’s known difficulties with mathematical calculation — after all, in order to know if the execution of the algorithm satisfies a property, it might have to first execute the algorithm, where GPT tends to make errors even with chain-of-thought prompting (Wei et al. 2022). However, our evidence suggests a deeper issue with categorizing inputs.
As an example, we asked the following question: “Give an example of a graph where the Ford-Fulkerson algorithm computes exactly six augmenting flows before terminating. Write the example as a list of edges and capacities.” This question is deliberately imprecise, and the intended answer is a graph with a source $s$, a sink $t$, and six intermediate vertices connected to both, leading to six parallel paths from s to t. However, more complex graphs could also be correct.
In the human survey, out of nine students, six described the graph made of parallel paths (the other three did not attempt the question). It is reasonable to expect that these students did not mentally execute the Ford-Fulkerson algorithm in order to verify their answer. Instead, they likely had some knowledge of the concept of parallel paths, and were able to leverage this knowledge to retrieve an example.
@@
Hedged Responses. Throughout GPT’s responses, we found frequent instances of GPT incorrectly ‘hedging’ its answers. When stating a true property of an algorithm, it frequently includes qualifiers such as ‘usually’ or ‘potentially’ that make the statement incorrect. E.g., responding to a question about the relationship between the Euclidean algorithm and the Fibonacci sequence, GPT-4 included the following line (emphasis ours):
@@
…Each Fibonacci number is the sum of the two preceding ones, with the sequence beginning as $F(0)=0$, $F(1)=1$, $F(2)=1$, $F(3)=2$, and so forth. This means each number in the sequence is relatively close to the sum of the two preceding numbers.
In another response, it stated that the golden ration $\phi$ was “one of the most irrational numbers”. In several responses, it stated that the Ford-Fulkerson algorithm ‘potentially’ tracks the capacity of reverse edges (not tracking would simply be an incorrect implementation of the algorithm). While this writing style may be an asset for subjects such as politics or health where being overly confident may cause harm, it causes statements about objective mathematical properties to be incorrect. This highlights the need for caution when using GPT-4 for teaching.
Hallucinations. All three versions of GPT occasionally produced hallucinations in the responses. The most common type of hallucination occurred when it tried to produce counter-examples. When asked to evaluate the (true) statement ‘GCD(a,b,c) = GCD(GCD(a,b), GCD(b,c)), GPT often tried to disprove it with a counter-example. GPT-3.5 generally claimed that its counter-example disproved the statement, despite the fact that the two sides of the equation were evidently the same.
The statement is false. Counterexample: Let a = 8, b = 12, c = 6. gcd(8, 12) = 4, gcd(12, 6) = 6. gcd(8, 12, 6) = 2, which is not equal to gcd(4, 6) = 2.
GPT-4 and GPT-4o both recognized when the ‘counter-examples’ failed to disprove the original statement. However, both continued to attempt to present counter-examples - in some cases, after many failed attempts, the responses became nonsensical. For example, from GPT-4:
…So GCD(GCD(1, 14))= 1). Therefore: GCD(6, 35, 14) = 1 And that verifies the consistency. Given the importance of a concept, premise restated true in a broader context, counter intuitive aligning confirm example specifics prove legitimacy …
Neither type of hallucination is observed in humans.
Discussion
We have presented a hierarchical scale for quantifying the understanding of algorithms. We verified its predictions empirically on human subjects and used it to compare generations of GPT and students. Our results show a significant improvement from GPT-3.5 to GPT-4/4o at all levels of algorithm understanding. We find that GPT-4 possesses a functional understanding of both the Euclidean algorithm and the Ford-Fulkerson algorithm, and it performs on a comparable level with CS PhD students. However, its reasoning abilities with mathematics lag behind its reasoning with language, and its understanding is not fully robust.
All versions of GPT were nearly perfect on code generation tasks. This is in line with other findings showing the competence of GPT in code generation (Vaithilingam, Zhang, and Glassman 2022; Savelka et al. 2023). This trend was not observed in the student respondents, who performed better at evaluating the algorithm than producing code on average. One possible reason for this is that the code for common algorithms, such as those tested, are prevalent in GPT’s training data. Code is highly structured, so even if the particular implementation of the algorithm has not been observed by GPT, it could replicate the changes, for example in variable names, by statistical inference.
Another trend is that GPT generally performed better on reasoning with language than with mathematics, while student performance was about the same. This difference goes beyond algebraic computations - GPT struggles with questions testing common-sense graph reasoning that humans can answer easily. However, for explanation questions and reasoning questions that do not involve examples, GPT-4 and 4o give consistently quality responses. This may suggest that mathematical examples play a larger role in human understanding than in LLMs.
@@ Discussion
Second, in order to determine whether GPT is reasoning, the concept of reasoning itself needs to be precisely defined. Our study indicates that GPT-4 has a sophisticated internal representation, e.g., its representation of the Euclidean algorithm includes its relationship to Linear Diophantine equations and mathematical properties of GCD. It is able to retrieve these properties under a variety of contexts. In a similar vein, it is likely that a student answering these questions has also been exposed to these properties. When evaluating the usefulness of a representation, we do not necessarily need to account for how the representation was created.
@@
Our results show that the hierarchy of understanding is consistent with classical notions of depth of understanding when tested on humans. While the results are also consistent with later versions of GPT having a ‘better’ understanding of the tested algorithms than undergraduates, such a conclusion does not follow. We worked with a limited population size, and the difference is confounded by other factors such as subject fatigue. Further research is needed to compare the quality of GPT and human responses to questions about algorithms. Despite these limitations, we feel that our scale makes progress towards a testable definition of understanding and can be extended to other algorithmic and similarly precise realms of understanding.
@@ Appendix A Survey Questions
Figure 7: The six point-scale shown to students to self-report their understanding, and the distribution of responses.
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common divisor of two integers a and b using the Euclidean algorithm.
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common divisor of two integers a and b using the Euclidean algorithm
@@
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common divisor of two integers a and b using the Euclidean algorithm.
@@
    You are speaking with a mathematics student who understands modular arithmetic. Explain the proof that the Euclidean algorithm finds the greatest common divisor of two numbers.